OpenStudy (anonymous):

A surveyor is standing 550 feet from the base of a redwood tree in the California Redwood Forest. The tree is 362 feet tall. Sketch a drawing that models the situation and then find the angle of elevation of the top of the tree from the spot where the surveyor is standing.

OpenStudy (anonymous):

 That's what I came up with and my answer is 33.4 degrees, am I right?

OpenStudy (anonymous):

\[\tan^{-1}(362/550)=0.582105\] \[0.582105\times(180/\pi)=33.3522\]

OpenStudy (anonymous):

@ZeHanz Does that look right to you?

OpenStudy (zehanz):

Fine! I set my calculator to degrees first, so I didn't have to do the second step...

OpenStudy (anonymous):

I would but I can't figure out how to do it on the Ti-nspire :(

OpenStudy (zehanz):

I've got one as well. Make a new document, choose 2:Add Graphs, then press menu. In the menu, press 9: Settings... Now you can set the units for angles.